Evaluate the expression
17! over 14! 3!

Found 2 solutions by Fombitz, Alan3354:
Answer by Fombitz(32388)
Answer by Alan3354(69443)

Evaluate the expression
17! over 14! 3!
======================
You can do it directly on a lot of calculators.
-->680
=============
Or,
17!/14! = 17*16*15 = 4080
4080/3! = 4080/6 = 680